What are the length and width of a rectanglar traffic sign if the length exceeds the width by 9 inches and the perimeter is 158 inches
Answer by rfer(16322)

2x+2(x+9)=158
2x+2x+18=158
4x=158-18
4x=140
x=140/4
x=35
x+9=44